# Guido Franchetti1 and Derek Harland2

### v1

@@ 1 Introduction
Moduli spaces and their metrics play a central role in the study of monopoles. The $L^{2}$ metric on the moduli space of monopoles in Euclidean space was introduced by Manton [23], who proposed that the geodesics of this metric should approximate the dynamics of slowly-moving monopoles. This conjecture was later confirmed by Stuart [28]. Inspired by Manton’s proposal, Atiyah–Hitchin calculated the metric on the moduli space of centred 2-monopoles [2]. Their calculation exploited the fact that the moduli space metrics are hyperkähler, and the Atiyah–Hitchin metric is now one of the best-known examples of a hyperkähler manifold. Many other prominent examples of hyperkähler metrics arise as moduli spaces of monopoles [22, 8, 9, 10]. The moduli space metric can be used to study not only classical but also quantum mechanics of monopoles [15]. Famously, Sen [25] made predictions for the $L^{2}$ cohomology of centred monopole moduli spaces, and these conjectures have been confirmed for charges 2 [18] and 3 [20].
@@
Monopoles on hyperbolic space share much mathematical structure with their Euclidean counterparts, including correspondences with spectral curves, rational maps, and (discrete) Nahm equations, so it would be natural to expect their moduli spaces to admit interesting metrics. However, it has been known for a long time that the natural $L^{2}$ metric on the hyperbolic monopole moduli space diverges [6]. In spite of this difficulty, many attempts have been made to discover natural geometric structures on their moduli spaces. Hitchin [17] constructed an Einstein metric on the moduli space of inversion-symmetric charge 2 monopoles using twistor theory. Unfortunately, it has proven difficult to generalise this approach to higher charges [19], and it is not clear what relevance this metric has to monopole dynamics. More recently, Bielawski–Schwachhofer [5, 4] introduced pluricomplex structures on the moduli spaces, building on twistor theory developed by Nash [24]. These structures also arose from an analysis of the supersymmetry of hyperbolic monopoles [11]. As with Hitchin’s work, the relevance of Bielawski–Schwachhofer’s approach to classical dynamics of monopoles is unclear because it is formulated in terms twistor theory.
@@
Gibbons–Warnick [16] and Franchetti–Ross [12] analysed the dynamics of well-separated monopoles, starting from a lagrangian for charged point particles. Interestingly, the metrics that they obtained were the LeBrun metrics, which had been studied much earlier in an entirely different context [21]. Franchetti–Schroers [13] calculated the $L^{2}$ metric for a circle-invariant 1-instanton on $S^{4}$, giving a finite metric on the moduli space of hyperbolic 1-monopoles which however is invariant only under a rotational subgroup of the $H^{3}$ isometry group. Finally, Sutcliffe [29] proposed a way to renormalise the divergent integral that defines the $L^{2}$ metric on the hyperbolic monopole moduli space. His boundary metric is related to the $L^{2}$ metric on the space of abelian connections on $S^{2}$, which had been considered earlier by Braam-Austin [7]. This metric was calculated explicitly in the case of inversion-symmetric charge 2 monopoles [30]. The asymptotic geometry of the boundary metric is very different from the metric of Franchetti–Ross [12, 14], so it is difficult to reconcile Sutcliffe’s metric with other approaches to monopole dynamics. It is also not clear that the boundary metric admits any analogue of the hyperkähler property.
@@
This article introduces a new approach to the geometry of hyperbolic monopoles. The starting point is a gauge-fixing condition introduced by Figueroa-O’Farrill–Gharamti [11]. On hyperbolic space with sectional curvature $-s^{2}$ it takes the form
@@
In this equation $a$ and $\phi$ are a Lie-algebra valued 1-form and 0-form describing a tangent vector to the moduli space. In the limit $s=0$ it reduces to the standard Coulomb gauge-fixing condition used to define the hyperkähler metric for Euclidean monopoles. This gauge-fixing condition was motivated by supersymmetry: it allowed the bosonic and fermionic zero-modes of the monopole to be related by supersymmetry transformations.
In this article we show that the gauge-fixing condition (1), subject to suitable boundary conditions, admits a unique solution. In other words, every tangent vector to the moduli space has a unique representative solving (1). We moreover show that the $L^{2}$ inner product of any two such tangent vectors is finite and so gives pointwise bilinear forms $g_{\pm}$ on the tangent bundle of the moduli space of hyperbolic monopoles. Thus the gauge fixing condition (1) solves the problem of the divergent $L^{2}$ metric.
The gauge-fixing condition (1) is unusual in that it explicitly involves the complex number $i$. So the solution $(a,\phi)$ should be expected to take values in a complex, rather than a real, Lie algebra. It is this fact that allows (1) to evade the divergence that afflicts more conventional approaches to $L^{2}$ metrics. However, this strength is also potentially a weakness: if $a$ and $\phi$ are not real then there is no reason to expect the integral of their square to be real. We nevertheless show that $g_{\pm}$ are real in at least two cases: the moduli space of charge 1 monopoles and the moduli space of inversion-symmetric charge $n$ monopoles. In particular, the hyperbolic analogue of the Atiyah–Hitchin metric is real.
We will moreover show that the gauge-fixing condition (1) leads to a natural decomposition
$T^{\mathbb{C}}M\cong E^{\pm}\otimes K^{\pm}$ (2)
of the complexified tangent bundle as a complex vector bundle $E^{\pm}$ and a 2-dimensional complex vector space $K^{\pm}$. This decomposition encodes the pluricomplex structures of Bielawski–Schwachhofer [5] and allows us to show that they are compatible with bilinear forms $g_{\pm}$. So $g_{\pm}$ appear to be natural hyperbolic counterparts to the hyperkähler metrics on moduli spaces of Euclidean monopoles. They also seem to be consistent with a physical interpretation of monopoles as particles whose inertial mass is equal to their energy. This is confirmed by an explicit calculation of metric components associated with Killing vector fields.
@@
Throughout this article we assume that the mass parameter $p\in\mathbb{R}_{>0}$ of the monopole is a half-integer. This has two benefits. First, it provides natural boundary conditions for (1) by identifying hyperbolic monopoles with circle-invariant instantons on $S^{4}$ [1]. And second, it allows for streamlined analytic proofs that exploit the compactness of $S^{4}$. Relaxing the assumption $2p\in\mathbb{N}$ would result in lengthier proofs involving more sophisticated analysis, which we leave for future work.
This article is organised as follows. Section 2 defines the hyperbolic monopole moduli spaces and states our main results for existence and uniqueness of solutions to (1) and finiteness of the $L^{2}$ integrals $g_{\pm}$, Theorems 3 and 4. Section 3 is devoted to the proof of Theorem 3. Section 4 establishes the decomposition (2) and explains the compatibility between $g_{\pm}$ and the pluricomplex structures. It also completes the proof of Theorem 4. Finally, section 5 contains some explicit calculations of $g_{\pm}$ in particular settings.
@@ 2.1 Preliminaries
A hyperbolic monopole is a pair $(A,\Phi)$ consisting of a connection $A$ and endomorphism $\Phi$ on a trivial vector bundle $V$ over hyperbolic space $H^{3}$ with sectional curvature $-1$. These are required to satisfy the Bogomolny equation,

### v2

@@ 1 Introduction
1.1 Background and motivation
An $SU(2)$ monopole consists of an $SU(2)$ connection $A$ and an $\mathfrak{su}(2)$-valued section $\Phi$ over a riemannian 3-manifold solving the BPS equation
$d^{A}\Phi=\ast F^{A}$ (1)
and suitable boundary conditions. The space of all solutions modulo gauge transformations is called the moduli space, and metrics on moduli spaces play a central role in the study of monopoles. The $L^{2}$ metric on the moduli space of monopoles in euclidean space is defined by the integral
$g=-\frac{1}{2}\int\operatorname{Tr}(a\wedge\ast a+\phi\wedge\ast\phi),$ (2)
in which $(a,\phi)$ are a Lie algebra-valued 1-form and 0-form which represent a tangent vector to the moduli space and which solve the linearisation of the BPS equation (1). A given tangent vector to the moduli space has many gauge-equivalent representatives $(a,\phi)$, so to make the metric (2) well-defined one must impose a gauge-fixing condition.
For monopoles on euclidean space, the gauge-fixing condition
$\ast d^{A}\ast a+[\Phi,\phi]=0$ (3)
admits a unique solution and leads to a well-defined metric on the moduli space. This metric was introduced by Manton [25], who proposed that the geodesics of this metric should approximate the dynamics of slowly-moving monopoles. This conjecture was later confirmed by Stuart [30].
The metric on the moduli space of euclidean monopoles is hyperkähler, and this fact allowed Atiyah–Hitchin to calculate the metric on the moduli space of centred 2-monopoles explicitly [2]. The Atiyah–Hitchin metric is now one of the best-known examples of a hyperkähler manifold. Many other prominent examples of hyperkähler metrics arise as moduli spaces of monopoles [24, 8, 9, 10].
The moduli space metric can be used to study not only classical but also quantum mechanics of monopoles [17]. Famously, Sen [27] made predictions for the $L^{2}$ cohomology of centred monopole moduli spaces, and these conjectures have been confirmed for charges 2 [20] and 3 [22].
Monopoles on hyperbolic space share much mathematical structure with their euclidean counterparts, including correspondences with spectral curves, rational maps, and (discrete) Nahm equations, so it would be natural to expect their moduli spaces to admit interesting metrics. However, it has been known for a long time that the gauge-fixing condition (3) leads to the integral (2) being divergent [6]. Thus the construction of an $L^{2}$-metric on the moduli space of hyperbolic monopoles is a long-standing open problem.
@@ 1.1 Background and motivation
In view of the difficulty with the $L^{2}$ metric, many alternative constructions of metrics and other geometric structures on hyperbolic monopole moduli spaces have been proposed. Hitchin [19] constructed an Einstein metric on the moduli space of inversion-symmetric charge 2 monopoles using twistor theory. Unfortunately, it has proven difficult to generalise this approach to higher charges [21], and it is not clear what relevance this metric has to monopole dynamics. More recently, Bielawski–Schwachhofer [5, 4] introduced pluricomplex structures on the moduli spaces, building on twistor theory developed by Nash [26]. Figueroa-O’Farrill–Gharamti [12] recovered these pluricomplex structures starting from an analysis of the supersymmetry of hyperbolic monopoles. As with Hitchin’s work, the relevance of Bielawski–Schwachhofer’s approach to the classical dynamics of monopoles is unclear.
@@
Gibbons–Warnick [18] and Franchetti–Ross [13] analysed the dynamics of well-separated monopoles, starting from a lagrangian for charged point particles. Interestingly, they showed that the dynamics were governed by the LeBrun metrics, which had been studied much earlier in an entirely different context [23]. Franchetti–Schroers [14] calculated the $L^{2}$ metric for a circle-invariant 1-instanton on $S^{4}$, obtaining a finite metric on the moduli space of hyperbolic 1-monopoles which is however invariant only under a rotational subgroup of the $H^{3}$ isometry group. Finally, Sutcliffe [31] proposed a way to renormalise the divergent integral that defines the $L^{2}$ metric on the hyperbolic monopole moduli space. His boundary metric, which had been considered earlier by Braam-Austin [7], is related to the $L^{2}$ metric on the space of abelian connections on $S^{2}$. This metric was calculated explicitly in the case of inversion-symmetric charge 2 monopoles [32]. The asymptotic geometry of the boundary metric is very different from the metric of Franchetti–Ross [13, 16], so it is difficult to reconcile Sutcliffe’s metric with other approaches to monopole dynamics. It is also not clear that the boundary metric admits any analogue of the hyperkähler property.
1.2 Main results
@@
This article introduces a new approach to the geometry of hyperbolic monopoles. The starting point is a gauge-fixing condition introduced by Figueroa-O’Farrill–Gharamti [12]. On hyperbolic space with sectional curvature $s^{2}$ it takes the form
@@
In this equation $a$ and $\phi$ are a Lie-algebra valued 1-form and 0-form describing a tangent vector to the moduli space. In the limit $s=0$ it reduces to the gauge-fixing condition (3) used to define the hyperkähler metric for euclidean monopoles. This gauge-fixing condition was motivated by supersymmetry: it allowed the bosonic and fermionic zero-modes of the monopole to be related by a supersymmetry transformation.
Our main result is
Theorem 1.
Let $(A,\Phi)$ be a $SU(2)$ hyperbolic monopole with half-integer mass. Then every tangent vector to the framed moduli space can be represented by a pair $(a_{\pm},\phi_{\pm})$ that solves (4). This representative is unique. Moreover, its $L^{2}$ integral (47) is finite and defines symmetric bilinear $\mathbb{C}$-valued pairings $g_{\pm}$ on the tangent space to the framed moduli space.
This result solves the problem of the divergent $L^{2}$ metric. We also prove an analogous result for the unframed moduli space, see Corollary 14 below. The assumption of half-integer mass means that $\|\Phi(x)\|$ approaches some constant $p\in\frac{1}{2}\mathbb{N}$ as $x$ approaches the boundary of hyperbolic space. The bilinear pairings $g_{\pm}$ in this theorem turn out to be complex conjugates of each other: $g_{-}=\overline{g_{+}}$.
The gauge-fixing condition (4) is unusual in that it explicitly involves the complex number $i$. As a consequence, the solution $(a,\phi)$ take values in the Lie algebra of the complexification of the gauge group. This is why the bilinear forms $g_{\pm}$ constructed in Theorem 1 are a priori $\mathbb{C}$-valued, rather than $\mathbb{R}$-valued. We nevertheless show that $g_{\pm}$ are real in two special situations.
The first situation is where we restrict to the moduli space of monopoles of charge 1. With our conventions this moduli space is diffeomorphic to $\mathbb{R}\times H^{3}$. In section 5 we find some solutions of the gauge-fixing condition (4), valid for any charge $n$ and any mass parameter $p$, and hence calculate $g_{\pm}(X,X)$ explicitly for particular choices of $X$. A special case of these calculations shows that, for $n=1$, $g_{\pm}$ is a constant real multiple of the product metric on $\mathbb{R}\times H^{3}$.
The second situation is where we restrict $g_{\pm}$ to the submanifold of the moduli space consisting of monopoles invariant under an inversion of hyperbolic space:
Proposition 2.
In particular, the hyperbolic analogue of the Atiyah–Hitchin metric is real.
Our second result concerns the structure of the tangent bundle of the moduli space:
Theorem 3.
The complexified tangent bundle $T^{\mathbb{C}}M^{f}_{n,p}$ of the framed moduli space $M^{f}_{n,p}$ of hyperbolic monopoles with mass parameter $p\in\frac{1}{2}\mathbb{N}$ and charge $n\in\mathbb{N}$ admits a natural decomposition
$T^{\mathbb{C}}M^{f}_{n,p}\cong E^{\pm}\otimes K^{\pm},$ (5)
in which $K^{\pm}$ is a complex 2-dimensional vector space and $E^{\pm}\to M^{f}_{n,p}$ is a rank $2n$ complex vector bundle.
The bundles $E^{\pm}$ in (5) are index bundles: their fibres are eigenspaces of a Dirac operator $D^{A}-\operatorname{ad}(\Phi)$ with eigenvalues $\pm\frac{i}{2}$. The vector spaces $K^{\pm}$ are the spaces of Killing spinors on $H^{3}$. There is a natural $L^{2}$ skew-symmetric pairing $\omega_{E^{\pm}}$ on the fibres of $E^{\pm}$, defined using the symplectic pairing on the spinor bundle of $H^{3}$, and also a natural symplectic pairing $\omega_{K^{\pm}}$ on $K^{\pm}$.
The decomposition (5) induces a family of almost complex structures parametrised by lines in $K^{\pm}$. For a given line $\langle\psi\rangle$ spanned by $\psi\in K^{\pm}$, the almost complex structure $J_{\langle\psi\rangle}$ is such that
$T^{1,0}_{\langle\psi\rangle}M^{f}_{n,p}=E^{+}\otimes\langle\psi\rangle.$ (6)
Our third main result concerns the compatibility of these structures with $g^{\pm}$:
Theorem 4.
The bilinear forms $g_{\pm}$, $\omega_{E^{\pm}}$, $\omega_{K^{\pm}}$ are related as follows:
$g_{\pm}=\frac{1}{2}\omega_{E^{\pm}}\otimes\omega_{K^{\pm}}.$ (7)
The bilinear form $g_{+}$ and almost complex structure $J_{\langle\psi\rangle}$ are compatible in the sense that
$g_{+}(X,X)=0\quad\forall X\in T^{1,0}_{\langle\psi\rangle}M^{f}_{n,p}.$ (8)
A decomposition similar to (5) was obtained by Nash [26] using less direct methods. Rather than working directly with monopoles, Nash worked with spectral curves, which are related to monopoles via a twistor corresondence. Nash constructed symplectic pairings on his analogues of $E^{\pm}$, while Bielawski–Schwachhofer [4, 5] studied the families of complex structures derived from Nash’s decomposition.
Conjecture 5.
The decomposition (5), skew-symmetric pairings $\omega_{E^{\pm}}$ and almost complex structures $J_{\langle\psi\rangle}$ coincide with those of Nash and Bielawski–Schwachhofer.
If true, this conjecture would imply that the $L^{2}$ metric $g_{\pm}$ is nondegenerate and that the almost complex structures $J_{\langle\psi\rangle}$ are integrable. But the conjecture is not straightforward to prove, because the passage from monopoles to spectral curves is complicated (we comment on this more in Appendix A).
@@
Throughout this article we assume that the mass parameter $p\in\mathbb{R}_{>0}$ of the monopole is a half-integer. This has two benefits. First, it provides natural boundary conditions for (4) by identifying hyperbolic monopoles with circle-invariant instantons on $S^{4}$, following [1]. Second, it allows for streamlined analytic proofs that exploit the compactness of $S^{4}$. Relaxing the assumption $2p\in\mathbb{N}$ would result in lengthier proofs which we leave for future work.
1.3 Outline
This article is organised as follows. Section 2 defines the hyperbolic monopole moduli spaces and gives a more precise statement of Theorem 1 (see Theorem 13). It also proves Proposition 2. The bulk of the proof of Theorem 1 is contained in section 3. Section 4 proves Theorems 3 and 4 (see Theorem 27 and Propositions 30 and 32). It also completes the proof of Theorem 1 in Proposition 28. Section 5 contains some explicit calculations of $g_{\pm}$ in particular settings. Finally, Appendix A compares our new definition of the framed moduli space with others that appear in the literature, and compares the almost complex structures $J_{\langle\psi\rangle}$ with the work of Bielawski–Schwachhofer [4, 5].
@@ 2.1 Preliminaries
A hyperbolic monopole is a pair $(A,\Phi)$ consisting of a connection $A$ and endomorphism $\Phi$ on a trivial vector bundle $V$ over hyperbolic space $H^{3}$ with sectional curvature 1-1. These are required to satisfy the Bogomolny equation,
@@
We will see shortly that hyperbolic monopoles with $2p\in\mathbb{N}$ correspond to circle-invariant instantons on $S^{4}$ [1]. This is due in part to the conformal equivalence $S^{1}\times H^{3}$ and $S^{4}\setminus S^{2}$. Let us construct this conformal equivalence explicitly, before explaining the relationship between instantons and monopoles.
We model $S^{4}$ as the set of unit vectors in $\mathbb{R}^{5}$. By writing $y_{1}=r_{1}\cos\theta$, $y_{2}=r_{1}\sin\theta$ with $r_{1}>0$, the metric on $\mathbb{R}^{5}\setminus\mathbb{R}^{3}$, where $\mathbb{R}^{3}\subset\mathbb{R}^{5}$ is defined by $r_{1}=0$, is seen to be conformally equivalent to the metric on $S^{1}\times H^{4}$:
The equation defining $S^{4}$, $r_{1}^{2}+y_{3}^{2}+y_{4}^{2}+y_{5}^{2}=1$, determines an embedding $H^{3}\subset H^{4}$. We thus obtain a conformal identification of $S^{1}\times H^{3}$ with $S^{4}\setminus S^{2}$. The action of the circle group on itself gives an action on $S^{1}\times H^{3}$ which in turn induces a rotation of $S^{4}$. The fixed set $S^{2}\subset S^{4}$ of the circle action is naturally identified with the boundary of $H^{3}$.
We now explain how circle-invariant instantons give rise to hyperbolic monopoles. Let $V\to S^{4}$ be a rank 2 bundle equipped with an $SU(2)$ lift of the following action of $\mathbb{R}$:
$t\cdot y=(y^{1}\cos(t)-y^{2}\sin(t),y^{1}\sin(t)+y^{2}\cos(t),y^{3},y^{4},y^{5}).$ (14)
Let $B$ be an instanton (that is, a connection with anti-self-dual curvature) on this bundle that is invariant under the action. We restrict $B$ to $S^{4}\setminus S^{2}$ and pull it back to $\mathbb{R}\times H^{3}$ via the map
$\mathbb{R}\times H^{3}\rightarrow S^{1}\times H^{3}\rightarrow S^{4}\setminus S^{2}$ (15)
where the first map is given by $(\theta,x)\mapsto(\mathrm{e}^{i\theta},x)$ and the second one is the conformal equivalence between $S^{1}\times H^{3}$ and $S^{4}\setminus S^{2}$ discussed previously. This gives a connection on $\mathbb{R}\times H^{3}$ that is invariant under the action of translations. After a gauge transformation, this connection takes the form $\Phi\,d\theta+A$, where $(A,\Phi)$ are a connection and section over $H^{3}$, pulled back to $\mathbb{R}\times H^{3}$. This connection has anti-self-dual curvature because the Hodge star is conformally invariant in middle dimension. The anti-self-dual equation for the curvature of this connection is equivalent to the Bogomolny equation (9) for $(A,\Phi)$. So $B$ determines a monopole provided that $(A,\Phi)$ satisfy the correct boundary conditions, which we verify in the next proposition.
Proposition 6.
Let $B$ be an irreducible anti-self-dual $SU(2)$ connection on a rank 2 bundle $V\to S^{4}$ that is invariant under an $SU(2)$ lift of the action (14). Then there exist $n,2p\in\mathbb{N}$ such that the following points hold.
  1. 1.
    The action of $2\pi\in\mathbb{R}$ on any fibre $V_{y}$ over $y\in S^{4}$ is multiplication by $(-1)^{2p}$.
  2. 2.
    The action of $t\in\mathbb{R}$ on any fibre $V_{y}$ over a fixed point $y$ has eigenvalues $e^{\mp ipt}$. The eigenbundles are line bundles $L^{\pm}$ over the fixed set $S^{2}$ such that $V|_{S^{2}}=L^{+}\oplus L^{-}$ and $c_{1}(L^{\pm})=\pm n$.
  3. 3.
    $B$ determines a hyperbolic monopole with charge $n$ and mass parameter $p$.
  4. 4.
    The second Chern number of $V$ is $2np$.
Proof.
Acting with $2\pi\in\mathbb{R}$ gives a section of $\mathrm{Aut}(V)$, because $2\pi$ fixes all points in $S^{4}$. This section is parallel because the $\mathbb{R}$-action fixes the connection $B$. Therefore its eigenvalues are constant. If it had two distinct eigenvalues then the holonomy group of $B$ would be a subgroup of $U(1)$. But this is not possible because $B$ is irreducible. So the eigenvalues are equal. Since the action is an $SU(2)$-action (meaning that it fixes a hermitian metric on $V$ and a non-vanishing section of $\Lambda^{2}V$), the determinant of the section of $\mathrm{Aut}(V)$ is 1. So the eigenvalues are $\pm 1$ and the section is $\pm\mathrm{Id}$.
Now consider the restriction of $V$ to the fixed set $S^{2}$. The action of $t\in\mathbb{R}$ gives an automorphism of $V|_{S^{2}}$. Again, the eigenvalues are constant and their product is 1. So they must be of the form $e^{\pm itp}$ for some $p\in\mathbb{R}$. Since $t=2\pi$ acts as $\pm 1$, the real number $p$ must satisfy $2p\in\mathbb{Z}$. Without loss of generality we can assume $p\geq 0$ and denote the eigenbundle for $e^{\pm itp}$ by $L^{\mp}$. Let $n=c_{1}(L^{+})\in\mathbb{Z}$. Since $c_{1}(V)=0$, $c_{1}(L^{-})=-n$. Since the connection $B$ is invariant, its restriction to $S^{2}$ is a sum of connections on $L^{\pm}$. The action of $t=2\pi$ on any fibre of $V|_{S^{2}}$ is given by multiplication with $(-1)^{2p}$, so by continuity, the same is true on any fibre of $V$.
Now we pull the bundle $V$ and connection $B$ back to $\mathbb{R}\times\overline{H}^{3}$, where $\overline{H}^{3}=H^{3}\cup S^{2}_{\infty}$. The bundle $V|_{S^{2}}$ is trivial, so the pulled back bundle over $\mathbb{R}\times S^{2}_{\infty}$ is also trivial. This trivialisation can be extended to a trivialisation of the bundle over $\mathbb{R}\times\overline{H}^{3}$. The $\mathbb{R}$-action on this trivial bundle takes the form
$t\cdot(\theta,x,v)=(\theta+t,x,h(t,\theta,x)v)\,\quad\forall(\theta,x,v)\in\mathbb{R}\times\overline{H}^{3}\times\mathbb{C}^{2},$ (16)
for some $h(t,\theta,x)\in SU(2)$. Our earlier calculation shows that $h(t,\theta,x)$ has eigenvalues $e^{\mp ipt}$ when $x\in S^{2}_{\infty}$. We can make a gauge transformation $g$ so that the $\mathbb{R}$-action is trivial. After doing so, the connection takes the form $\Phi\,d\theta+A$. So we have that
$B$ $=g\Phi g^{-1}\,d\theta+gAg^{-1}-dgg^{-1}$ (17)
$h(t,\theta,x)$ $=g(\theta+t,x)g(\theta,x)^{-1}.$ (18)
When $x\in S^{2}_{\infty}$, the contraction of $B$ with $\partial/\partial\theta$ is zero, because $B$ was pulled back from a connection that extends across $S^{2}\subset S^{4}$. So the equations together imply that
$g\Phi g^{-1}=\frac{\partial g}{\partial\theta}g^{-1}=\frac{\partial}{\partial t}\bigg{|}_{t=0}h(t,\theta,x)$ (19)
for $x\in S^{2}_{\infty}$. Therefore the eigenvalues of $\Phi$ over $S^{2}_{\infty}$ are the same as the eigenvalues of $\partial_{t}h$ over $S^{2}\subset S^{4}$. Since $h$ has eigenvalues $e^{\mp ipt}$ over $S^{2}$, the eigenvalues of $\Phi$ are $\mp ipt$, and the eigenspaces are identified with the eigenbundles $L^{\pm}$ of $h$.
By Chern-Weil theory and (12),
$c_{2}(V)=\frac{1}{8\pi^{2}}\int_{[0,2\pi]\times H^{3}}\operatorname{Tr}(F^{B}\wedge F^{B})=\frac{-1}{4\pi^{2}}\int_{[0,2\pi]\times H^{3}}d\theta\wedge\operatorname{Tr}(d^{A}\Phi\wedge F^{A})\\ =\frac{-1}{2\pi}\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge F^{A})=2np.$ (20)
Now we describe the reverse process. Let $(A,\Phi)$ be a hyperbolic monopole; we will attempt to construct a circle-invariant instanton. We choose a gauge transformation $g:\mathbb{R}\times H^{3}\to SU(2)$ and define a connection $B$ on $\mathbb{R}\times H^{3}$ via (17). We want this connection to be periodic in $\theta$; for this it is sufficient to choose $g$ such that
$g(\theta+2\pi,x)=\pm g(\theta,x).$ (21)
With this choice, $B$ descends to an anti-self-dual connection on $S^{1}\times H^{3}\simeq S^{4}\setminus S^{2}$. We now attempt to extend this connection to $S^{4}$. The holonomy of the connection along $S^{1}\times\{x\}$ is
$g(2\pi,x)^{-1}\exp(-2\pi\Phi(x))g(0,x).$ (22)
As $x$ approaches $S^{2}_{\infty}$ the eigenvalues of this matrix tend to $\pm e^{2\pi ip}$ and $\pm e^{-2\pi ip}$. A theorem of Sibner–Sibner shows that the connection extends to $S^{4}$ if and only if the eigenvalues approach 1 [28, 29]. In the “++” case this means that $p$ is an integer, and in the “-” case it means that $p\in(\frac{1}{2}\mathbb{Z})\setminus\mathbb{Z}$. Thus:
Proposition 7.
A hyperbolic monopole extends to a circle-invariant instanton on $S^{4}$ if and only if $2p\in\mathbb{N}$.
2.2 The framed and unframed moduli spaces
The Bogomolny equation is invariant under gauge transformations, and it is natural to study the moduli space of solutions modulo gauge transformations. In fact, there are two natural moduli spaces to consider, the framed and unframed moduli spaces. We denote by $C_{n,p}$ the space of solutions of the Bogomolny equations and boundary conditions with charge $n$ and mass parameter $p$ that extend smoothly to the bundle $V\to S^{4}$. We define the group of unframed gauge transformations to be the group $G^{u}$ of functions $g:H^{3}\to SU(2)$ that extend to smooth circle-invariant sections of the bundle $\operatorname{Aut}(V)\to S^{4}$. Note that circle-invariance demands that on the fixed $S^{2}\subset S^{4}$, $g$ maps $L^{+}$ to $L^{+}$ and $L^{-}$ to $L^{-}$. Two monopoles are called unframed gauge equivalent if they are related by an unframed gauge transformation. The moduli space $M^{u}_{n,p}$ of unframed monopoles is the quotient of $C_{n,p}/G^{u}$. This is known to be a smooth manifold of dimension 4n14n-1 [7].
To define the framed moduli space, we need to recall the definition of the Chern–Simons number. Let $A^{0},A^{1}$ be two connections on $H^{3}$ and let $\tilde{A}=(1-t)A^{0}+tA^{1}$ be a connection on $0,1]\times H^{3}$. The Chern-Simons number of $A^{0},A^{1}$ is defined to be